Problem asml SNF 2009-03-05 17:40:07: Non Criticals Only

gsosa at snf.stanford.edu gsosa at snf.stanford.edu
Fri Mar 6 16:22:39 PST 2009


 Vendors cleaned the E-table and tested with hotspot wafer- Looks good. Vendors checked and adjusted white light focus/tilt., performed red to white light tilt matching and red/blue test. Vendors also did BMC calibrations and edge sensor/mark sensor calibrations. New system reference state defined. IQC tests look good. System is up and ready to use. Please expose and inspect a test wafer if possible to ensure your work. 




More information about the asml-pcs mailing list